# BERTs are Generative In-Context Learners

### v1

@@ 4 Evaluation
4.1 Language understanding (SuperGLUE)
[…]
Results
We show the resulting scores of evaluation with the same prompts as GPT-3 in Table 1. DeBERTa clearly outperforms its contemporary and scales much more favorably than the family of GPT models (Figure 1). Interestingly, the average performance of the 1.4B DeBERTa gets close to the reported performance of the largest 175B GPT-3 ($68.4$ vs. $68.9$, 1-shot). However, this average score is still far from the performance of a finetuned DeBERTa, which is more than 20 percentage points higher (He et al., 2021); the average few-shot performance of DeBERTa is slightly better than a finetuned BERT-large (Devlin et al., 2019; Wang et al., 2019).
@@ Results
We show the in-context-learning results from this group in Table 2, where we evaluate the 1.4B DeBERTa with a comparable GPT-3 model. The scores showcase consistently stronger performance of the masked language model, similarly to the language understanding tasks. One difference to those tasks is the rate of scaling, which appears to be similar between the two types of language models (Figure 1).
Table 2: Results of text completion, language modeling and Winograd-style tasks. All tasks are measured with accuracy, we show the performance of the largest available DeBERTa (1.4 billion parameters) and of a similarly-sized GPT-3 model, the best results are boldfaced.
@@
The SacreBLEU scores on each language pair are given in Table 3. Unlike in the previous two task groups, the tables have turned and the causal language model clearly outperforms the masked model in all comparisons. We believe that the subpar performance of DeBERTa can be (at least) in part explained by its relatively small and clean monolingual training corpus (Section 3), because the performance on this task is highly dependent on the presence of multilingual data in the corpus (Lin et al., 2022). The rate of improved translation performance with larger scale appears to be similar between the two models (Figure 1).
Table 3: Machine translation results. We report SacreBLEU scores (Post, 2018) with signature BLEU+case.mixed+numrefs.1+smooth.exp+tok.intl+version.1.2.20 (higher is better). The table shows the performance of the largest available DeBERTa (1.4 billion parameters) and of a similarly-sized GPT-3 model, the best results are boldfaced.
@@
Table 4: Closed-book question answering and commonsense reasoning. The first three tasks are measured with the exact-match accuracy and the rest is measured with classification accuracy. The table shows the performance of the largest available DeBERTa (1.4 billion parameters) and of a similarly-sized GPT-3 model, the best results are boldfaced.
@@
4.4 Closed-book question answering and commonsense reasoning
Results
The question-answering performance is given in Table 4. Apparently, the results of DeBERTa are substantially worse on closed-book question answering, compared to GPT-3. We believe that this highlights a more general disadvantage of the MLM training objective – the model can often retrieve world knowledge from the rich bidirectional context during training, not needing to store it in its learned weights; similar effect has been shown in retrieval-augmented language models (Samuel et al., 2024). However, the commonsense reasoning abilities are comparable between the two models. The scaling behavior is again similar between the two models (Figure 1). The same is also true about the improvement when given more in-context examples (Figure 4), which are especially important for the tasks evaluated with exact-match accuracy, where the goal is not only to answer correctly but also to match the expected style and form of the gold answers.
@@ 5 Related work
Our study is not the first to propose few-shot learning for pretrained BERT-like models, but to our best knowledge, we are the first to demonstrate in-context learning abilities of these models. The major difference is that previous studies rely on few-shot finetuning, adapting the model to a downstream task by updating its weights. Studies by Schick and Schütze (2021), Gao et al. (2021), or Xia et al. (2022) demonstrated that few-shot finetuning is a promising method. Unfortunately, these methods can be cumbersome to implement in practice and the simplicity of in-context learning is often preferred even though its performance is generally found to be worse (Liu et al., 2022).
Scaling up masked language models
We selected DeBERTa as the largest publicly available masked language model that can handle large context size. However, we should note that there are even larger BERT-like models: Megatron BERT with 3.9 billion parameters by Shoeybi et al. (2019, it is unfortunately unavailable and without a mechanism for length generalization), and XLM-RoBERTa with 10.7 billion parameters by Goyal et al. (2021, it is unfortunately multilingual and also limited to sequences of 512 tokens).
@@
We believe that an important take-away from our study is that there is an opportunity for merging the observed strengths of masked language models with the strengths of causal language models. There have already been several attempts to create such a hybrid model: T5 (Raffel et al., 2020) and GLM (Du et al., 2022) proposed autoregressive fill-in-the-blank training objectives, CM3 is based on a causal-mask objective (Aghajanyan et al., 2022), and PrefixLM uses a partially bidirectional causal language model (Raffel et al., 2020). In particular, Ding et al. (2024) convincingly argues that prefix language models are better suited for in-context learning than causal language models.
@@ 6 Conclusion
In this paper, we empirically demonstrated that masked language models can be capable in-context learners. These models, often assumed to be deprecated and only useful for finetuning, are shown to be competitive with causal language models on in-context learning: with similar absolute performance, similar scaling behavior, and with similar improvement when given more few-shot demonstrations.
Although the DeBERTa language model was not designed to be used for (generative) in-context learning, we proposed inference methods that enable it to be used in this setting without any additional training (or any other change of its learned weights). We argue that in a future work, its results could be easily improved by pretraining on a larger and more diverse text corpus, by scaling up the parameter count and by pretraining with a longer context length. Moreover, there is even more potential in merging the strength of causal language modeling with masked language modeling, instead of relying only on one of these training objectives.
@@
I would like to thank Lilja Øvrelid, Andrey Kutuzov, and Erik Velldal for providing valuable feedback, and for their never-ending encouragement and support.
@@
Appendix A Evaluation details
@@
Few-shot prompting.
For each evaluated sample, the example demonstrations are randomly selected (without replacement) from the training set of each task; if the training set is not available, we sample from the only available dataset split, making sure not to select the same sample as the evaluated one. We format these examples using the respective prompt templates and concatenate them together, joined by two newline characters. The numbers of shots used for each task are given in Appendix D.
A.1 Needle in a haystack
@@ A.4 Translation
A.5 Closed-book question answering
@@
Appendix B Examples of text generation
To also give a sense of the quality of the text produced by DeBERTa, we include some examples of text generation in this section. We use the exact same example prompts that were used in the GPT-3 paper (Brown et al., 2020), to provide a fair estimate of the generative qualities. All text completions were generated with nucleus sampling (Holtzman et al., 2020). [^5] Using these hyperparameters: top_k=64, top_p=0.9, temperature=0.2. We compare the largest DeBERTa 1.4B with OPT 1.3B, an openly available replication of the GPT-3 1.3B (Zhang et al., 2022).
B.1 Learning and using novel words
Results
{minted}
{minted}
{minted}
{minted}
{minted}
{minted}
B.2 Correcting English grammar
Results
{minted}
{minted}
{minted}
{minted}
{minted}
{minted}
{minted}
Appendix C Ablation study of ranking implementation
{minted}
Results
In Table 5, we compare the original implementation of PLL by Salazar et al. (2020) that only masks the target subword; our approach that also masks next two subwords; other two alternatives that mask two or four subwords in total; and the exact unidirectional computation of log-likelihood (using the same input formatting as for generation). The additional masks clearly help to make better estimates while the exact computation seems to not be appropriate for inherently bidirectional models.
Table 5: Ablation study of different ranking methods applied to DeBERTa. Evaluated using zero-shot setting and the largest DeBERTa 1.4B model on ReCoRD.
Ranking method ReCoRD (EM) ReCoRD (F1)
Pseudo log-likelihood; 1 mask (Salazar et al., 2020) 80.9 81.6
Pseudo log-likelihood; 2 masks 86.0 86.8
Pseudo log-likelihood; 3 masks (our method) 87.1 87.9
Pseudo log-likelihood; 4 masks 86.9 87.8
Exact log-likelihood 77.2 77.8
Appendix D All results
For reference, we list all results of the DeBERTa models evaluated throughout this paper in Table 6. The GPT-3 results that were used for comparison are published in Brown et al. (2020).
Table 6: Results for all evaluations performed in this paper. The second and third column shots the dataset splits and evaluation metrics, both of them replicating the GPT-3 evaluation setup. Note the BLEU scores used for evaluating translation are SacreBLEU scores with signature BLEU+case.mixed+numrefs.1+smooth.exp+tok.intl+version.1.2.20.
Task Split Metric $n$ shots 0-shot (1.4B) 1-shot (0.1B) 1-shot (0.4B) 1-shot (0.9B) 1-shot (1.4B) $n$-shot (1.4B)

### v2

@@ Results
We show the in-context-learning results from this group in Table 2, where we evaluate the 1.5B DeBERTa with a comparable GPT-3 model. The scores showcase consistently stronger performance of the masked language model, similarly to the language understanding tasks. One difference to those tasks is the rate of scaling, which appears to be similar between the two types of language models (Figure 1).
Table 2: Results of text completion, language modeling and Winograd-style tasks   All tasks are measured with accuracy, we show the performance of the largest available DeBERTa (1.4 billion parameters) and of a similarly-sized GPT-3 model, the best results are boldfaced.
@@
The SacreBLEU scores on each language pair are given in Table 3. Unlike in the previous two task groups, the tables have turned, and the causal language model clearly outperforms the masked model in all comparisons. We believe that the subpar performance of DeBERTa can be (at least) in part explained by its relatively small and clean monolingual training corpus (Section 3), because the performance on this task is highly dependent on the presence of multilingual data in the corpus (Lin et al., 2022). The rate of improved translation performance with larger scale appears to be similar between the two models (Figure 1).
Table 3: Machine translation results   We report SacreBLEU scores (Post, 2018) with signature BLEU+case.mixed+numrefs.1+smooth.exp+tok.intl+version.1.2.20 (higher is better). The table shows the performance of the largest available DeBERTa (1.4 billion parameters) and of a similarly-sized GPT-3 model, the best results are boldfaced.
@@
4.4 Closed-book question answering and commonsense reasoning
Results
The question-answering performance is given in Table 4. Apparently, the results of DeBERTa are substantially worse on closed-book question answering compared to GPT-3. We believe that this highlights a more general disadvantage of the MLM training objective – the model can often retrieve world knowledge from the rich bidirectional context during training, not needing to store it in its learned weights; similar effect has been shown in retrieval-augmented language models (Samuel et al., 2024). However, the commonsense reasoning abilities are comparable between the two models. The scaling behavior is again similar between the two models (Figure 1). The same is also true about the improvement when given more in-context examples, which are especially important for the tasks evaluated with exact-match accuracy, where the goal is not only to answer correctly but also to match the expected style and form of the gold answers (Figure 4).
Table 4: Closed-book question answering and commonsense reasoning   The first three tasks are measured with the exact-match accuracy and the rest is measured with classification accuracy. The table shows the performance of the largest available DeBERTa (1.4 billion parameters) and of a similarly-sized GPT-3 model, the best results are boldfaced. A detailed description of the evaluation method is given in Appendix E, full results are in Appendix F.
@@ 5 Related work
While our work demonstrates the emergence of in-context learning in masked language models, prior research has explored different approaches to few-shot learning with these architectures. The dominant paradigm has been few-shot finetuning, where the model’s weights are updated using a small number of examples. Studies by Schick and Schütze (2021), Gao et al. (2021), and Xia et al. (2022) showed promising results with this approach. However, these methods require additional training steps with a complicated training objective, making them more complex to implement compared to the simple prompting-based in-context learning demonstrated in our work. Despite the generally lower performance of in-context learning compared to few-shot finetuning (Liu et al., 2022), its simplicity and immediacy have made it the preferred choice in many practical applications.
Other large masked language models
Our choice of DeBERTa for this study was motivated by its unique combination of size and capability to handle extended context lengths. While larger masked language models exist, such as Megatron BERT with 3.9 billion parameters (unfortunately not publicly available; Shoeybi et al., 2019) and XLM-RoBERTa with 10.7 billion parameters (Goyal et al., 2021), they have limitations that make them less suitable for studying in-context learning. Megatron BERT lacks mechanisms for length generalization, which is crucial for processing long prompts with multiple examples, while XLM-RoBERTa’s multilingual nature and restricted sequence length of 512 tokens would confound our analysis. DeBERTa’s architecture, particularly its relative positional embeddings, makes it an ideal candidate for exploring how masked language models scale with in-context learning.
@@
Our empirical findings, particularly the complementary strengths of masked and causal models demonstrated in Section 4, suggest significant potential in combining these approaches. Several architectures have already explored this direction, even if inadvertently: T5 (Raffel et al., 2020), BART (Lewis et al., 2020) and GLM (Du et al., 2022) introduced autoregressive fill-in-the-blank objectives; CM3 developed a causal-mask approach (Aghajanyan et al., 2022); and PrefixLM implemented a partially bidirectional causal model (Dong et al., 2019; Raffel et al., 2020). These efforts align with our observations about the distinct advantages of masked and causal objectives. The recent work by Ding et al. (2024) provides theoretical support for this direction, demonstrating that prefix language models, which combine aspects of both architectures, are particularly well-suited for in-context learning.
@@ 6 Conclusion
This paper demonstrates that masked language models can be capable in-context learners. We show that these models – often considered deprecated and limited only to finetuning – can match and sometimes even exceed the performance of their causal counterparts in this domain. Our evaluation reveals that masked and causal models exhibit remarkably similar characteristics in terms of overall performance, scaling behavior, and improvements with additional in-context demonstrations. Most notably, we validate these capabilities using DeBERTa without any architectural modifications or additional training. We achieve this through carefully designed inference methods that unlock the model’s latent generative abilities.
Our findings point to several promising directions for future research. First, DeBERTa’s performance could likely be enhanced through straightforward improvements such as training on larger and more diverse corpora, increasing model scale, and extending the pretraining context length. More fundamentally, the complementary strengths we observed between masked and causal models – where each architecture excels in different tasks – suggest an exciting opportunity to develop hybrid approaches that combine the best of both paradigms. Rather than viewing these as competing architectures, future work might explore how to synthesize their distinct advantages into more capable and versatile language models.
These results argue for a broader reconsideration of how we approach language model architecture and training. The field’s recent focus on causal models, while productive, may have prematurely discounted the potential of alternative approaches that are not limited to unidirectional text processing. Our work demonstrates that the path forward likely involves embracing architectural diversity rather than converging on a single dominant paradigm.
@@
I am deeply grateful to Lilja Øvrelid, Andrey Kutuzov, and Erik Velldal for providing insightful feedback, for their never-ending encouragement and support, and for making Oslo a warm and welcoming place.
@@
Appendix A Examples of text generation
To also give a sense of the quality of the text produced by DeBERTa, we include some examples of text generation in this section. We use the exact same example prompts that were used in the GPT-3 paper (Brown et al., 2020), to provide a fair estimate of the generative qualities. All text completions were generated with nucleus sampling (Holtzman et al., 2020). [^5] Using these hyperparameters: top_k=64, top_p=0.9, temperature=0.2. We compare the largest DeBERTa 1.5B with OPT 1.3B, an openly available replication of the GPT-3 1.3B (Zhang et al., 2022).
A.1 Learning and using novel words
Results
{minted}
{minted}
{minted}
{minted}
{minted}
{minted}
A.2 Correcting English grammar
Results
{minted}
{minted}
{minted}
{minted}
{minted}
{minted}
{minted}
Appendix B Ablation study of text generation
We empirically evaluate different approaches for text generation with DeBERTa language models. In particular, we study how many additional mask tokens should be used during autoregressive generation, and we also compare our approach with previously proposed methods based on Markov-chain Monte Carlo sampling (Wang and Cho, 2019). For this analysis, we use German-to-English machine translation as a representative generative task. We evaluate different generation methods using one-shot prompts (following Section 2.1) and the largest 1.5B DeBERTa model. The translation quality is measured with SacreBLEU score, using the same signature as in the main experiments.
Results
The results shown in Table 5 demonstrate that using additional mask tokens during generation substantially improves the performance of DeBERTa. This aligns with our hypothesis that additional masks help to reduce the effect of the end-of-sequence token on the generated text. The results also show that while adding a fourth mask token still marginally improves the performance, the gain is negligible compared to using three mask tokens.
We also compare our autoregressive approach with the Gibbs sampling method proposed by Wang and Cho (2019). There we relied on the same hyperparameters that are suggested in the official implementation: 500 total iterations, 250 burn-in iterations with top-100 sampling and temperature of $1.0$. [^6] Available on GitHub: https://github.com/nyu-dl/bert-gen. Their sampling-based approach performs substantially worse than autoregressive generation, regardless of the initialization strategy. We noticed that it often produces infinite token repetitions or locally-coherent but globally-disconnected pieces of text.
Table 5: Ablation study of different generation methods applied to DeBERTa   Evaluated using one-shot setting and the largest DeBERTa 1.5B model on German-to-English translation with SacreBLEU score.
[…]
Autoregressive generation; 1 mask 10.0
Autoregressive generation; 2 masks 22.4
Autoregressive generation; 3 masks (our method) 23.7
Autoregressive generation; 4 masks 23.9
Markov-chain Monte-Carlo (random initialization; Wang and Cho, 2019) 1.6
Markov-chain Monte-Carlo (mask initialization; Wang and Cho, 2019) 2.6
Appendix C Ablation study of ranking implementation
{minted}
Previous work
After running the initial experiments, we were informed about the related work by Kauf and Ivanova (2023). Their study addresses a similar problem of the naive pseudo-log-likelihood scoring as this paper, but they do not target strongly correlated contiguous tokens in general, they focus on words that are split into multiple tokens – as they observed that PLL overestimates the likelihood of those sequences. We include their two proposed solutions in the comparison of different ranking methods.
Results
In Table 6, we compare the original implementation of PLL by Salazar et al. (2020) that only masks the target subword; our approach that also masks next two subwords; other two alternatives that mask two or four subwords in total; the PLL-word-l2r and PLL-whole-word scoring functions by Kauf and Ivanova (2023); and the exact unidirectional computation of log-likelihood (using the same input formatting as for generation). The additional masks clearly help to make better estimates while the exact computation seems to not be appropriate for inherently bidirectional models.
Table 6: Ablation study of different ranking methods applied to DeBERTa   Evaluated using zero-shot setting and the largest DeBERTa 1.5B model on ReCoRD.
Ranking method ReCoRD (EM) ReCoRD (F1)
Pseudo log-likelihood; 1 mask (Salazar et al., 2020) 80.9 81.6
Pseudo log-likelihood; 2 masks 86.0 86.8
Pseudo log-likelihood; 3 masks (our method) 87.1 87.9
Pseudo log-likelihood; 4 masks 86.9 87.8
Pseudo log-likelihood; word–l2r (Kauf and Ivanova, 2023) 78.2 78.8
Pseudo log-likelihood; whole–word (Kauf and Ivanova, 2023) 75.8 76.5
Exact log-likelihood (unidirectional) 77.2 77.8
Appendix D Detailed SuperGLUE results
This appendix provides a detailed analysis of DeBERTa’s performance on the SuperGLUE benchmark, focusing on two aspects: the relationship between the number of demonstration examples and model performance (Figure 5), and the task-specific scaling behavior (Figure 6).
Effect of shot count
Figure 5 demonstrates how DeBERTa’s performance changes as we increase the number of in-context examples from 0 to 32. Unlike the main results where we optimize the number of shots for each subtask independently, here we deliberately use the same number of shots across all tasks to provide a controlled comparison with GPT-3’s results from Brown et al. (2020); there, SuperGLUE tasks are the only ones with such detailed few-shot evaluation.
The plot reveals several interesting patterns:
  1. 1.
    The steepest improvement occurs between 0 and 4 shots, suggesting that even a small number of examples provides substantial benefit.
  2. 2.
    The overall trend is very similar between DeBERTa and GPT-3, which again shows that masked language models can be just as good in-context learners as causal language models.
  3. 3.
    The performance of DeBERTa decreases after 8 or more shots. We believe that this is mainly caused by its imperfect processing of contexts longer than 512 tokens – as discussed in Section 2.3.
Figure 5: The average performance on the SuperGLUE benchmarks as a function of number of shots     As opposed to the other SuperGLUE few-shot results, where we select the number of shots for each subtask according to the performance on its training split, here all subtasks are evaluated with the same number of shots. In this way, we can compare DeBERTa 1.5B directly to Figure 3.8 from Brown et al. (2020), which gives the same evaluation for GPT 175B (unfortunately not for smaller, more comparable, models).
Task-specific scaling analysis
Figure 6 breaks down the scaling behavior for each SuperGLUE task individually, revealing the varying impacts of model size across different language understanding challenges. The plots demonstrate that while DeBERTa and GPT-3 both exhibit generally positive scaling trends, their scaling patterns differ substantially across task types. Most notably, DeBERTa shows steeper scaling curves than GPT-3 on the majority of tasks. This suggests that bidirectional attention mechanisms may provide particular advantages for scaling some language understanding capabilities.
Figure 6: Detailed evaluation of natural language understanding   As opposed to the overall results presented in Figure 1, these plots show the scaling behavior on each subtask of the SuperGLUE benchmark. We can see the DeBERTa out-scales GPT-3 consistently across all tasks, with the exception of WiC where both models fail to show any learning ability.
Appendix E Evaluation details
@@
Few-shot prompting
For each evaluated sample, the example demonstrations are randomly selected (without replacement) from the training set of each task; if the training set is not available, we sample from the only available dataset split, making sure not to select the same sample as the evaluated one. We format these examples using the respective prompt templates and concatenate them together, joined by two newline characters. The numbers of shots used for each task are given in Appendix F.
E.1 Needle in a haystack
@@ E.4 Translation
E.5 Closed-book question answering
@@
Appendix F All results
For reference, we list all results of the DeBERTa models evaluated throughout this paper in Table 7. The GPT-3 results that were used for comparison are published in Brown et al. (2020, Table H.1).
Table 7: Results of all evaluations performed in this paper   The second and third column shots the dataset splits and evaluation metrics, both of them replicating the GPT-3 evaluation setup. Note the BLEU scores used for evaluating translation are SacreBLEU scores with signature BLEU+case.mixed+numrefs.1+smooth.exp+tok.intl+version.1.2.20.
Task Split Metric $n$ shots 0-shot (1.5B) 1-shot (0.1B) 1-shot (0.4B) 1-shot (0.9B) 1-shot (1.5B) $n$-shot (1.5B)